OpenStudy (anonymous):

Solve for x: 2(x+8)^1/2 - 5(x-13)^1/2

OpenStudy (anonymous):

is the equation =0?

OpenStudy (anonymous):

it doesn't say it on the page. But I'm assuming yes.

OpenStudy (anonymous):

If the equaltion =0 (2x+16)^1/2=(5x-65)1/2 square both sides you have 2x+16=5x-65 3x=81 x=27

OpenStudy (anonymous):

how'd you get 3x?

OpenStudy (anonymous):

-2x both side

OpenStudy (anonymous):

Right! Okay, thank you.